OpenStudy (anonymous):

3log(x+4)-2log(x-7)+5log(x-2)-log(x^2) ..how do I write this as one logarithm?

OpenStudy (ghazi):

\[\log(x+4)^3-\log(x-7)^2+\log(x-2)^5-logx^2=\log[ \frac{ (x+4)^2(x-2)^5 }{ (x-7)^2x^2 }]\]

OpenStudy (ghazi):

use these properties \[\log a + \log b = \log(ab)-----\log a- \log b = \log \frac{ a }{ b }\] hope that helps you :)